Find the probability that Z <=   0.3283
Find the probability that Z <=   1.4008
Find the probability that Z <=   1.2458
Find the probability that Z >=0.9973
Find the probability that Z >=0.2499
Find the probability that Z >=-0.0433
Find the probability that -0.4455   <=   Z   <=   -0.0874
Find the probability that -0.3063   <=   Z   <=   0.3892
Find the probability that -0.9015   <=   Z   <=   0.3261
Answer #1

Answer)

a) p(z<0.3283) = 0.6293 (from z table round off 0.3283 to 0.33 then observe 0.3 from the column and 0.03 from the row, the value is 0.6293)

similarly

b) p(z<1.4008) = 0.9192

c) p(z<1.2458) = p(z<1.25) = 0.8944

d) p(z>0.9973) = 1-p(z<0.9973) = 1-p(z<1.00) = 1-0.8413 = 0.1587

e) p(z>2499) = 1-p(z<25) = 1-0.8944 = 0.1056

f) p(z>-0.0433) = 1-p(z<-0.04) = 1-0.4840 = 0.5160

g) p(-0.4455   <=   Z   <=   -0.0874) = p(z<-0.0874) - p(z<-0.4455)

= p(z<-0.09) - p(z<-0.45)

= 0.4641 - 0.330

= 0.1341

h) p( -0.3063   <=   Z   <=   0.3892)

= p(z<0.39)-p(z<-0.31)

= 0.6517 - 0.3783

= 0.2734

i) p(-0.9015   <=   Z   <=   0.3261)

= p(z<0.33) - p(z<0.-90)

= 0.6293-0.1841

= 0.4452